# references.bib

### v1

@@ Abstract
For a graph family $\mathcal{F}$, let $\mathrm{ex}(n,\mathcal{F})$ and $\mathrm{spex}(n,\mathcal{F})$ denote the maximum number of edges and maximum spectral radius of an $n$-vertex $\mathcal{F}$-free graph, respectively, and let $\mathrm{EX}(n,\mathcal{F})$ and $\mathrm{SPEX}(n,\mathcal{F})$ denote the corresponding sets of extremal graphs. Wang, Kang, and Xue showed that if $\mathrm{EX}(n,\mathcal{F})$ consists of Turán graphs $T_{n,r}$ plus $O(1)$ edges, then $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$ for $n$ large enough. Fang, Tait, and Zhai extended this result by showing if $e(T_{n,r})\leq\mathrm{ex}(n,\mathcal{F})<e(T_{n,r})+\lfloor n/2r\rfloor$ then $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$ for $n$ large enough. In this paper we extend the result further and in many cases we can show that our result is best possible, answering a question of Fang, Tait, and Zhai.
@@ 1 Introduction
In this paper we ask for a condition on the value of $\mathrm{ex}(n,\mathcal{F})$ which implies some relation between $\mathrm{EX}(n,\mathcal{F})$ and $\mathrm{SPEX}(n,\mathcal{F})$. In [byrne2024general] this question was addressed when $\mathcal{F}$ contains a bipartite graph, so here we focus on the case where all graphs in $\mathcal{F}$ are non-bipartite. The first positive result in this direction was proved by Wang, Kang, and Xue [wang2023conjecture], answering a conjecture of Cioabă, Desai, and Tait [cioabua2022spectral]. Recently, the result of [wang2023conjecture] was extended by Fang, Tait, and Zhai [fang2024turan] (here $\chi(\mathcal{F})$ denotes $\min\{\chi(F):F\in\mathcal{F})\}$).
@@
Thus, Theorem 1 shows that $c(r)\geq 1/2r$ for every $r$. In this paper we give improved lower and upper bounds on $c(r)$ when $r+1\geq 4$ and in many cases we can determine $c(r)$ exactly (see 1 below). When $r=2$, we prove that $c(2)\leq 1/3$ and conjecture that this is the correct value. These results also improve the upper bound $c(2)\leq 3/2$ obtained by considering the odd wheel $W_{15}$ [yuan2020extremal, cioabua2022spectral] (we are not aware of existing upper bounds for larger $r$).
@@
We answer this question in the negative for all $r$ (see the proofs of 1 and 2).
@@ 1.1 Notation and definitions
We write $K_{n}$ for the complete graph on $n$ vertices; $K_{V_{1},\ldots,V_{r}}$ for the complete multipartite graph with partite sets $V_{1},\ldots,V_{r}$; $T_{n,r}$ for the $r$-partite Turán graph on $n$ vertices; $P_{\ell}$ for a path on $\ell$ vertices; and $M_{n}$ for a matching on $n$ vertices (or almost matching if $n$ is odd). For a graph family $\mathcal{F}$, let $\chi(\mathcal{F})=\min\{\chi(F):F\in\mathcal{F}\}$. Given a graph $G$, we write $V(G)$ and $E(G)$ for its vertex set and edge set, respectively; given a set $S\subseteq V(G)$, we write $N_{G}(S):=\{v\in V(G):v\sim u\text{ for all }u\in S\}$, or just $N(S)$ when the underlying graph is understood; $G[U]$ for the induced subgraph on $U$, $E[U]$ for $E(G[U])$ and $e(U)$ for $|E(U)|$; $G[U,W]$ for the bipartite subgraph induced by $U$ and $W$; $A(G)$ for the adjacency matrix of $G$ and $\lambda(G)$ for the spectral radius of $A(G)$. For a vertex $v\in V(G)$ and a set $S\subseteq V(G)$, we write $N_{S}(v):=N(v)\cap S$ and $d_{S}(u)$ for $|N_{S}(u)|$; we write $G-v$ for the graph obtained by deleting the vertex $v$; and $d^{\prime}(v)$ for the number of walks of length two starting from $v$. For $n\in\mathbb{N}$, let $U(G,n)$ be a graph on $n$ vertices consisting of as many disjoint copies of $G$ as possible with isolated vertices in the remainder. We write $G\cup H$ and $G+H$ for the disjoint union and join, respectively, of $G$ and $H$ and $k\cdot G$ for the union of $k$ disjoint copies of $G$.
@@ 1.2 Main results
Our main result extends Theorem 1 in the case $r\geq 3$. Define
$c_{1}(r):=\frac{1}{r}\left(1-\left[\frac{r-1}{2}\left(2+\frac{5}{r-1}-\frac{4}{r}+\sqrt{\left(2+\frac{5}{r-1}-\frac{4}{r}\right)^{2}-\frac{4}{r-1}\left(\frac{6}{r-1}-\frac{4}{r}\right)}\right)\right]^{-1}\right)$
@@
Then there is a finite family of graphs $\mathcal{F}$ such that $\chi(F)=r+1$, $\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,r})+\frac{k-1}{kr}n+O(1)$, and $\mathrm{SPEX}(n,\mathcal{F})\cap\mathrm{EX}(n,\mathcal{F})=\emptyset$ for an infinite sequence of $n$.
@@
We may conclude from Theorem 2 and 1 that $c(r)\sim 1/r$ as $r\to\infty$.
@@ 1.3 Applications
We are not aware of any other existing extremal results in this category. However, we are interested in constructing small graphs $F$ for which we can show $\mathrm{ex}(n,F)=e(T_{n,r})+\frac{n}{2r}+O(1)$. Other than the path power discussed above, the simplest such $F$ we found is pictured below.
@@ 2 Proof of Theorem 2
Our overall strategy is as follows. In subsection 2.1 we prove a general result about the structure of the graphs in $\mathrm{EX}(n,\mathcal{F})$. In subsection 2.2, we use this structural information to prove Theorem 2.
@@
Given a graph family $\mathcal{F}$ with $\chi(F)=r+1\geq 3$ and $\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,r})+Qn$ where $Q<1/r$, let $k=k(\mathcal{F})$ be the maximum integer such that for some tree $T_{k}$ of order $k$, for $m$ large enough, $U(T_{k},m)+T_{m(r-1),r-1}$ is $\mathcal{F}$-free.
Since $Q<1/r$, $k$ is finite; more precisely, if $n=mr$ then $(n/r+O(1))\frac{k-1}{k}\leq Qn$, hence $k\leq\lfloor 1/(1-Qr)\rfloor$.
@@ 2.1 Structure of edge extremal graphs
Suppose $\mathcal{F}$ is a finite family of graphs, $\chi(\mathcal{F})=r+1\geq 3$, and for $n$ large enough, any graph obtained by adding more than $Qnr$ edges to $T_{nr,r}$ is not $\mathcal{F}$-free, where $Q<1/r$. Then for $n$ large enough, $\mathrm{EX}(n,\mathcal{F})$ contains a graph $G$ obtained from $T_{n,r}$ by embedding trees of some fixed order $k$ in one part, then making a bounded number of edge deletions and additions to the whole graph.
@@
  • Let $m^{\prime}$ be the sum of the orders of the $C_{i}$, then $m^{\prime}\geq m$ and $(C_{1}\cup\cdots\cup C_{m})+T_{m(r-1),r-1}$ is obtained from $T_{m^{\prime}r,r}$ by adding $m^{\prime}$ edges to one part, which contradicts $Q<1/r$.
Assume that $G=G_{n}$ is any sequence of graphs in $\mathrm{EX}(n,\mathcal{F})$. Let $m$ be a large enough even constant satisfying 2. Let $M=\max\{|V(F)|:F\in\mathcal{F}\}$. Choose constants $\varepsilon,\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}$ satisfying the following inequalities (we do not attempt to avoid redundancy).
@@
$2|X|\sqrt{\varepsilon}n\leq\sum_{v\in X}d_{B}(x)=2e(B)\leq\varepsilon n^{2}\Longrightarrow|X|\leq\sqrt{\varepsilon}n.$
For $v\in V_{i}$, we say that $v$ is ordinary if $N(v)=\bigcup_{j\neq i}V_{j}-X$.
@@
Suppose not. Note than an ordinary vertex $u$ satisfies $d(u)\geq(1-1/r-\sqrt{\varepsilon})n+O(1)$.
@@
If $G^{\prime}$ is obtained from $G$ by deleting some vertex and adding an ordinary vertex $u\in V_{i}$, then $G^{\prime}$ is $\mathcal{F}$-free.
@@
If we delete $v$ and add an ordinary vertex we obtain
@@
and if we delete $v$ and add an ordinary vertex we obtain
@@
Now that we have redefined the partition of $V$, we redefine an ordinary vertex $v\in W_{i}$ to mean one such that $N(v)=V-W_{i}$.
@@
If $d_{W_{j}}(v)\leq w_{j}-m-1$ then by 7 replacing $v$ with an ordinary vertex increases the number of edges, and by 6 and $G\in\mathrm{EX}(n,\mathcal{F}$) this is a contradiction. ∎
@@
Assume for a contradiction that $e(W_{1}),e(W_{2})$ are both arbitrarily large. By 7, $G[W_{1}],G[W_{2}]$ have bounded degree, so this implies we can find a matching $M_{1}$ in $W_{1}$ of order $m$ and a matching $M_{2}$ in $W_{2}$ of order $2(m+1)m$. For each $v\in M_{1}$, by 8, there are at most $m$ edges in $M_{2}$ either of whose endpoints are not adjacent to $v$, which implies that there are some $(m+1)m-m^{2}=m$ edges in $M_{2}$ inducing a matching $M_{2}^{\prime}$ such that $G[M_{1},M_{2}^{\prime}]\simeq K_{m,m}$. Similarly to lemmas proved above, having found $v_{31},\ldots,v_{im}$, we have
@@
For each $i\neq 1$, all vertices in $W_{i}^{o}$ are ordinary.
@@
We can write $W_{1}=W_{1}^{o}\sqcup W_{1}^{m}$, where $G[W_{1}^{o}]$ is a disjoint union of trees of order $k$, each tree component of $G[W_{1}^{o}]$ occurs $\omega(1)$ times, and $|W_{1}^{m}|=O(1)$.
@@
The graph $G[W_{1}]$ has at most $m$ components whose order is at least $k+1$.
@@
The graph $G[W_{1}]$ has at most $m$ disjoint cycles of order at most $k$.
@@
The graph $G[W_{1}]$ has at least $\sqrt{n}$ tree components of order $k$.
@@
Assume for a contradiction then $G^{\prime}\supseteq F\in\mathcal{F}$. Then $S^{\prime}:=V(F)\cap S\neq\emptyset$. Note there are at least $\sqrt{n}/2$ copies of $T^{\prime}$ which do not intersect $F$ in any vertex. Now we have
@@
$\sum_{i=1}^{a}|V(T_{i})|\leq(k-1)^{2}k.$
@@
Combining 3, 4, and 7 gives the result (we put any trees of order $k$ which only occur $O(1)$ times into $W_{1}^{m}$). ∎
@@
Every vertex in $W_{1}^{o}$ is ordinary.
@@
Assume for a contradiction that $v\in W_{1}^{o}$ and $N(v)\not\supseteq V-W_{1}$, so there is some $u\not\in W_{1}$ with $v\not\sim u$. Then the graph $G^{\prime}$ obtained by adding the edge $uv$ contains some $F\in\mathcal{F}$. Let the component of $G[W_{1}]$ containing $v$ be $T$. By 11, there are $\omega(1)$ disjoint copies of $T$ in $G[W_{1}]$, $\omega(1)$ of which do not intersect $F$ in any vertex. Copying the proof of 6, we find that $N(N_{F}(v))$ contains such a copy of $T$ so that $F\subseteq G$, a contradiction. ∎
At this point we know that $G$ can be obtained from $K_{W_{1},\ldots,W_{r}}$ by embedding trees of order $k$ in $W_{1}$, then deleting and adding $O(1)$ edges.
@@
Suppose $H$ is any graph obtained from $K_{W_{1},\ldots,W_{R}}$ by embedding trees in $W_{1}$, then deleting and adding $O(1)$ edges. Let $W_{i}^{o},W_{i}^{m}$ be defined as above, and assume that $|W_{i}^{o}|\to\infty$, $|W_{i}^{m}|=O(1)$ for each $i$. For any $i,j$, let $H_{ij}$ be defined by deleting a vertex $u$ in $W_{i}^{o}$ such that $d_{W_{i}}(u)\leq 1$ and adding a vertex $v$ whose neighborhood is $V-W_{j}-u$. If $H$ is $\mathcal{F}$-free then so is $H_{ij}$. Moreover, $e(H_{ij})-e(G)=|W_{i}|-|W_{j}|-1$ if $i\neq 1$ and $e(H_{ij})-e(G)\geq|W_{i}|-|W_{j}|-2$ if $i=1$.
@@
By the above, we are done unless $w_{1}=w_{i}+2$ for some $i$. Then by 13, $G_{1i}$ is $\mathcal{F}$-free and $e(G_{1i})-e(G)\geq 0$. Performing this operation decreases $\sum_{i<j}|w_{i}-w_{j}|$. Thus by performing this operation successively, we eventually obtain an $\mathcal{F}$-free graph $G^{\prime}$ which satisfies $|w_{i}-w_{j}|\leq 1$ for every $i,j$ and $e(G^{\prime})\geq e(G)$, hence $G^{\prime}\in\mathrm{EX}(n,\mathcal{F})$. ∎

### v2

@@ Abstract
For a graph family $\mathcal{F}$, let $\mathrm{ex}(n,\mathcal{F})$ and $\mathrm{spex}(n,\mathcal{F})$ denote the maximum number of edges and maximum spectral radius of an $n$-vertex $\mathcal{F}$-free graph, respectively, and let $\mathrm{EX}(n,\mathcal{F})$ and $\mathrm{SPEX}(n,\mathcal{F})$ denote the corresponding sets of extremal graphs. Wang, Kang, and Xue showed that if $\mathrm{EX}(n,\mathcal{F})$ consists of Turán graphs $T_{n,r}$ plus $O(1)$ edges, then $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$ for $n$ large enough. Fang, Tait, and Zhai extended this result by showing if $e(T_{n,r})\leq\mathrm{ex}(n,\mathcal{F})<e(T_{n,r})+\lfloor n/2r\rfloor$ then $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$ for $n$ large enough. In this paper we extend the result further and for many values of $r$ we can show that our result is best possible, answering a question of Fang, Tait, and Zhai.
@@ 1 Introduction
In this paper we ask for a condition on the value of $\mathrm{ex}(n,\mathcal{F})$ which implies some relation between $\mathrm{EX}(n,\mathcal{F})$ and $\mathrm{SPEX}(n,\mathcal{F})$. In [byrne2024general] this question was addressed when $\mathcal{F}$ contains a bipartite graph, so here we focus on the case where all graphs in $\mathcal{F}$ are non-bipartite. The first positive result in this direction was proved by Wang, Kang, and Xue [wang2023conjecture], answering a conjecture of Cioabă, Desai, and Tait [cioabua2022spectral]. Recently, the result of [wang2023conjecture] was extended by Fang, Tait, and Zhai [fang2024turan]. Below, $\chi(\mathcal{F})$ denotes $\min\{\chi(F):F\in\mathcal{F})\}$.
@@
Thus, Theorem 1 shows that $c(r)\geq 1/2r$ for every $r$. In this paper we give improved lower and upper bounds on $c(r)$ when $r\geq 3$ and in many cases we can determine $c(r)$ exactly (see 1 below). When $r=2$, we prove that $c(2)\leq 1/3$ and conjecture that this is the correct value. These results also improve the upper bound $c(2)\leq 3/2$ obtained by considering the odd wheel $W_{15}$ [yuan2020extremal, cioabua2022spectral] (we are not aware of existing upper bounds for larger $r$).
@@
We answer this question in the negative for all $r\geq 3$ (see the proof of 1).
@@ 1.1 Notation and definitions
We write $K_{n}$ for the complete graph on $n$ vertices; $K_{V_{1},\ldots,V_{r}}$ for the complete multipartite graph with partite sets $V_{1},\ldots,V_{r}$; $T_{n,r}$ for the $r$-partite Turán graph on $n$ vertices; $P_{\ell}$ for a path on $\ell$ vertices; and $M_{n}$ for a matching on $n$ vertices (or almost matching if $n$ is odd). For a graph family $\mathcal{F}$, let $\chi(\mathcal{F})=\min\{\chi(F):F\in\mathcal{F}\}$. Given a graph $G$, we write $V(G)$ and $E(G)$ for its vertex set and edge set, respectively; given a set $S\subseteq V(G)$, we write $N_{G}(S):=\{v\in V(G):v\sim u\text{ for all }u\in S\}$, or just $N(S)$ when the underlying graph is understood; $G[U]$ for the induced subgraph on $U$, $E[U]$ for $E(G[U])$ and $e(U)$ for $|E(U)|$; $G[U,W]$ for the bipartite subgraph induced by $U$ and $W$; $A(G)$ for the adjacency matrix of $G$ and $\lambda(G)$ for the spectral radius of $A(G)$. For a vertex $v\in V(G)$ and a set $S\subseteq V(G)$, we write $N_{S}(v):=N(v)\cap S$ and $d_{S}(v)$ for $|N_{S}(v)|$; we write $G-v$ for the graph obtained by deleting the vertex $v$; and $d^{\prime}(v)$ for the number of walks of length two starting from $v$. For $n\in\mathbb{N}$, let $U(G,n)$ be a graph on $n$ vertices consisting of as many disjoint copies of $G$ as possible with isolated vertices in the remainder. We write $G\cup H$ and $G+H$ for the disjoint union and join, respectively, of $G$ and $H$ and $k\cdot G$ for the union of $k$ disjoint copies of $G$.
@@ 1.2 Main results
Our main result extends Theorem 1 in the case $r\geq 3$. Let $\alpha$ be the largest root of the quadratic
$f(x)=-\frac{1}{r-1}x^{2}+\left(2+\frac{5}{r-1}-\frac{4}{r}\right)x+\left(\frac{4}{r}-\frac{6}{r-1}\right),$
and let $c_{1}(r)=\frac{\alpha-1}{\alpha r},$ i.e.
$c_{1}(r)=\frac{1}{r}\left(1-\left[\frac{r-1}{2}\left(2+\frac{5}{r-1}-\frac{4}{r}+\sqrt{\left(2+\frac{5}{r-1}-\frac{4}{r}\right)^{2}-\frac{4}{r-1}\left(\frac{6}{r-1}-\frac{4}{r}\right)}\right)\right]^{-1}\right).$
@@
The appearance of the constant $c_{1}(r)$ is explained as follows. The ‘worst case scenario’ in the proof of Theorem 2 occurs when an edge extremal graph $H$ is obtained from $T_{n,r}$ by embedding paths $P_{k}$ in a large part along with one extra edge, while a spectral extremal graph $G$ is obtained from $T_{n,r}$ by embedding stars $K_{1,k-1}$ in a small part. (The $k$-vertex trees with smallest and largest spectral radius are $P_{k}$ and $K_{1,k-1}$, respectively, so intuitively they should also give the extremal contributions to the spectral radius of a larger graph in which they are embedded.) We will show that
$\lambda(H)-\lambda(G)=\left[2-\frac{k-5+6/k}{r-1}-\frac{4(k-1)}{kr}\right]\frac{1}{n}+O(n^{-2}).$
The first term inside the brackets accounts for the extra edge; the second term comes from the difference between embedding stars versus paths; and the last term comes from the difference between embedding the trees in a small part versus a large part. In order to avoid this potential counterexample to Theorem 2, we require that the bracketed expression be positive, which is equivalent to $k<\alpha$. The number of edges embedded in the large part of $H$ is $\frac{k-1}{kr}n+O(1)<c_{1}(r)n$. Regarding the other possibility $k>\alpha$, we can always construct an $\mathcal{F}$ which gives rise to this ‘worst case scenario,’ which proves that the constant $c_{1}(r)$ in Theorem 2 is sharp up to some integrality considerations:
@@
Then there is a finite family of graphs $\mathcal{F}$ such that $\chi(F)=r+1$, $\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,r})+\frac{k-1}{kr}n+O(1)$, and $\mathrm{SPEX}(n,\mathcal{F})\cap\mathrm{EX}(n,\mathcal{F})=\emptyset$ for an infinite sequence of $n$. Thus, $c(r)\leq\frac{k-1}{kr}.$
@@
Additionally, the asymptotics of $c(r)$ can be deduced from Theorem 2 and 1, namely $c(r)\sim 1/r$ as $r\to\infty$.
@@ 1.3 Applications
We are not aware of any other existing extremal results in this category. However, with the goal of constructing graphs $F$ for which we can show $\mathrm{ex}(n,F)=e(T_{n,r})+\frac{n}{2r}+O(1)$ we began exploring specific small graphs. Other than the path power discussed above, the simplest such $F$ we found is described as follows. The graph $F_{1}$ consists of two paths on four vertices $bcde$, $fghi$, with additional edges $bf,bg,cf,ch,dg,di,eh,ei$ and additionally one dominating vertex $a$.
@@ 2 Proof of Theorem 2
Our overall strategy is as follows. In subsection 2.1 we prove 1, which provides general structural information about the graphs in $\mathrm{EX}(n,\mathcal{F})$. In subsection 2.2, we prove that the same general structural description also applies to graphs in $\mathrm{SPEX}(n,\mathcal{F})$. The difference in spectral radius of two graphs $H,G$ satisfying this structural description is equal to $2(e(H)-e(G))/n+O(n^{-2})$, plus an error term which depends on what trees are embedded in $H$ and $G$ and in which part of their $r$-partition. We show that this error term is at most $\left[-\frac{k-5+6/k}{r-1}-\frac{4(k-1)}{kr}\right]\frac{1}{n}+O(n^{-2})$. If $e(G)=e(H)$ then $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$, while otherwise $e(G)<e(H)$ and we have
$\lambda(H)-\lambda(G)\geq\left[2-\frac{k-5+6/k}{r-1}-\frac{4(k-1)}{kr}\right]\frac{1}{n}+O(n^{-2}).$
The assumption $Q<c_{1}(r)$ in Theorem 2 guarantees that the bracketed term is positive, giving a contradiction.
@@
Given a graph family $\mathcal{F}$ with $\chi(F)=r+1\geq 3$ and $\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,r})+Qn$ for large enough $n$ where $Q<1/r$, let $k=k(\mathcal{F})$ be the maximum integer such that for some tree $T_{k}$ of order $k$, for all $m$, $U(T_{k},m)+T_{m(r-1),r-1}$ is $\mathcal{F}$-free.
Since $Q<1/r$, $k$ is finite; more precisely, whenever $k|m$ we have $m\frac{k-1}{k}\leq Qmr$, hence $k\leq\lfloor 1/(1-Qr)\rfloor$.
@@ 2.1 Structure of edge extremal graphs
Suppose $\mathcal{F}$ is a finite family of graphs, $\chi(\mathcal{F})=r+1\geq 3$, and for $n$ large enough, any graph obtained by adding more than $Qnr$ edges to $T_{nr,r}$ is not $\mathcal{F}$-free, where $Q<1/r$. Then for $n$ large enough, $\mathrm{EX}(n,\mathcal{F})$ contains a graph $G$ obtained from $T_{n,r}$ by embedding disjoint trees of some fixed order $k$ in one part, then making a bounded number of edge deletions and additions to the whole graph.
@@
  • Let $m^{\prime}$ be the sum of the orders of the $C_{i}$, then $m^{\prime}\geq m$ and $(C_{1}\cup\cdots\cup C_{m})+T_{m^{\prime}(r-1),r-1}$ is obtained from $T_{m^{\prime}r,r}$ by adding $m^{\prime}$ edges to one part, which contradicts $Q<1/r$.
Assume that $G=G_{n}$ is any sequence of graphs in $\mathrm{EX}(n,\mathcal{F})$. Let $m$ be a large enough even constant satisfying 2. Let $M=\max\{|V(F)|:F\in\mathcal{F}\}$. Choose constants $\varepsilon,\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}$ satisfying the following inequalities (we do not attempt to avoid redundancy or optimize the constants).
@@
$2|X|\sqrt{\varepsilon}n\leq\sum_{v\in X}d_{B}(v)=2e(B)\leq 2\varepsilon n^{2}\Longrightarrow|X|\leq\sqrt{\varepsilon}n.$
For $v\in V_{i}$, we say that $v$ is $V$-ordinary if $N(v)=\bigcup_{j\neq i}V_{j}-X$.
@@
Suppose not. Note than a $V$-ordinary vertex $u$ satisfies $d(u)\geq(1-1/r-\sqrt{\varepsilon})n+O(1)$.
@@
If $G^{\prime}$ is obtained from $G$ by deleting some vertex and adding a $V$-ordinary vertex $u\in V_{i}$, then $G^{\prime}$ is $\mathcal{F}$-free.
@@
If we delete $v$ and add a $V$-ordinary vertex we obtain
@@
and if we delete $v$ and add an $V$-ordinary vertex we obtain
@@
Call a vertex $v\in W_{i}$ $W$-ordinary if $N(v)=V-W_{i}$.
@@
If $d_{W_{j}}(v)\leq w_{j}-m-1$ then by 7 replacing $v$ with a $W$-ordinary vertex increases the number of edges, and by 6 and $G\in\mathrm{EX}(n,\mathcal{F}$) this is a contradiction. ∎
@@
Assume for a contradiction that $e(W_{1}),e(W_{2})$ are both arbitrarily large. By 7, $G[W_{1}],G[W_{2}]$ have bounded maximum degree, so this implies we can find a matching $M_{1}$ in $W_{1}$ of order $m$ and a matching $M_{2}$ in $W_{2}$ of order $2(m+1)m$. For each $v\in M_{1}$, by 8, there are at most $m$ edges in $M_{2}$ either of whose endpoints are not adjacent to $v$, which implies that there are some $(m+1)m-m^{2}=m$ edges in $M_{2}$ inducing a matching $M_{2}^{\prime}$ such that $G[M_{1},M_{2}^{\prime}]\simeq K_{m,m}$. Similarly to lemmas proved above, having found $v_{31},\ldots,v_{im}$, we have
@@
For each $i\neq 1$, all vertices in $W_{i}^{o}$ are $W$-ordinary.
@@
We can write $W_{1}=W_{1}^{o}\sqcup W_{1}^{m}$, where $G[W_{1}^{o}]$ is a disjoint union of trees of order $k$, each tree component of $G[W_{1}^{o}]$ occurs more than $M(m+1)$ times, and $|W_{1}^{m}|=O(1)$.
@@
The number of components of $G[W_{1}]$ whose order is at least $k+1$ is at most $m$.
@@
The graph $G[W_{1}]$ contains at most $m$ disjoint cycles of order at most $k$.
@@
The graph $G[W_{1}]$ has at least $\sqrt{n}$ isomorphic tree components of order $k$.
@@
Assume for a contradiction that $G^{\prime}\supseteq F\in\mathcal{F}$. Then $S^{\prime}:=V(F)\cap S\neq\emptyset$. Note there are at least $\sqrt{n}/2$ copies of $T^{\prime}$ which do not intersect $F$ in any vertex. Now we have
@@
If the tree components of $G[W_{1}]$ of order at most $k-1$ are $T_{1},\ldots,T_{j}$, then we have
$\sum_{i=1}^{j}|V(T_{i})|\leq(k-1)^{2}k.$
@@
Combining 3, 4, and 7 gives the result (we put any trees of order $k$ which occur at most $M(m+1)$ times in $W_{1}^{m}$). ∎
@@
Every vertex in $W_{1}^{o}$ is $W$-ordinary.
@@
Assume for a contradiction that $v\in W_{1}^{o}$ and $N(v)\not\supseteq V-W_{1}$, so there is some $u\not\in W_{1}$ with $v\not\sim u$. Then the graph $G^{\prime}$ obtained by adding the edge $uv$ contains some $F\in\mathcal{F}$. Let the component of $G[W_{1}]$ containing $v$ be $T$. By 11, there are more than $M(m+1)$ disjoint copies of $T$ in $G[W_{1}]$, and more than $Mm$ of these do not intersect $F$ in any vertex. Then from 8 we have
$|N(N_{F}(v))\cap W_{1}|\geq w_{1}-Mm$
so some copy of $T$ is contained in $N(N_{F}(v))$. This implies $F\subseteq G$, a contradiction. ∎
At this point we know that $G$ can be obtained from $K_{W_{1},\ldots,W_{r}}$ by embedding disjoint trees of order $k$ in $W_{1}$, then deleting and adding $O(1)$ edges.
@@
Suppose $H$ is any graph obtained from $K_{W_{1},\ldots,W_{r}}$ by embedding disjoint trees in $W_{1}$, then deleting and adding $O(1)$ edges. Let $W_{i}^{o},W_{i}^{m}$ be defined as above, and assume that $|W_{i}^{o}|\to\infty$, $|W_{i}^{m}|=O(1)$ for each $i$. For any $i,j$, let $H_{ij}$ be defined by deleting a vertex $u$ in $W_{i}^{o}$ such that $d_{W_{i}}(u)\leq 1$ and adding a vertex $v$ whose neighborhood is $V-W_{j}-u$. If $H$ is $\mathcal{F}$-free then so is $H_{ij}$. Moreover, $e(H_{ij})-e(G)=|W_{i}|-|W_{j}|-1$ if $i\neq 1$ and $e(H_{ij})-e(G)\geq|W_{i}|-|W_{j}|-2$ if $i=1$.
@@
By the above, we are done unless $w_{1}=w_{i}+2$ for some $i$. Then by 13, $G_{1i}$ is $\mathcal{F}$-free and $e(G_{1i})-e(G)\geq 0$, i.e. $G_{1i}\in\mathrm{EX}(n,\mathcal{F})$. One can check that $G_{1i}$ satisfies $|w_{i}-w_{j}|\leq 1$ for all $i,j$. ∎